# Quantum Entanglement Distribution via Uplink Satellite Channels

### v2

@@ I Introduction
We first introduce our uplink setup and the various physical phenomena affecting the Bell measurement, such as beam widening, beam wandering, atmospheric turbulence and attenuation, stray photons, and mode-mismatch. We ignore polarization error [24, 25] as it is negligible. We also ignore the Doppler effect [26, 27] as it can be compensated. Then, we combine all these phenomena to derive expressions for Fidelity and Efficiency of the protocol. Next, we compare these metrics against various important parameters, for which we use a realistic range of values. Lastly, we discuss our results and potential future directions towards the realization of uplink satellite networks. A detailed derivation of the Fidelity and Efficiency expressions can be found in the supplementary material, which, along with the Mathematica code containing the simulations, can be found here: https://shorturl.at/UtHxy.
@@ II The Protocol
Figure 1: A schematic of the the proposed uplink setup. The setup consists of two ground stations sending one-half of their Bell pairs to a satellite via an uplink channel. The Bell measurement happens on the satellite, entangling the other two ground photons (represented by the dark blue curvy line). Inset: the Bell-measurement apparatus inside the satellite. The apparatus consists of two optical modes, directed by a Polarizing Beam Splitter (PBS), followed by two 45 polarizers (one on each mode), and further followed by two PBSes, one on each mode, and foue bucket detectors, one on each output of the final PBSes.
@@ IX Results and discussion
Figure 2: Fidelity $F$ and success probability $\eta_{tot}$ vs time-gating window for various wavepacket widths $\sigma_{t}$ (in nm). The satellite altitude is taken to be 500km and the ground stations are separated by a distance of 1000km. All other parameters are chosen according to Table 1 of the Supplementary Material
@@
Results and Analysis. For our analysis, we fix the satellite equidistant from the two ground stations. To study how success probability (8) and fidelity (9) behave, we consider realistic parameter values chosen from various experiments and industry standards [see Table 1 and Section 9 of the supplementary material]. We organize our simulation around four main quantities of interest: temporal wavepacket width $\sigma_{t}$, time-gating window $t=t_{\mathrm{max}}-t_{\mathrm{min}}$, satellite altitude $h$ and distance $D_{G}$ between ground stations. The first two quantities have a large range of experimentally realizable values and affect the overall fidelity and $\eta_{tot}$ non-trivially, therefore their trends must be examined to establish a reasonable range of values for these parameters. Following this, we characterize the range of distances and satellite altitudes across which the protocol can produce feasible results.
@@
Best performance. Optimising the design of entanglement distribution in the framework of our protocol is a challenging task because the problem is a nonlinear optimization. Nevertheless, the most performant protocol we can reach is with a gating window of 40 ns, wavepacket width of 10nm, satellite altitude of 200km, and ground separation of 300km: this protocol distributes entangled state at a fidelity of 0.972 with an success probability of $1.5\times 10^{-4}$. Nonetheless, even with a satellite altitude of 500km and a ground separation as far as 1000km, we can obtain fidelity of 0.84 and a success probability of $2.404\times 10^{-6}$.

### v3

@@ I Introduction
We first introduce our uplink setup and the various physical phenomena affecting the Bell measurement, such as beam widening, beam wandering, atmospheric turbulence and attenuation, stray photons, and mode-mismatch. We ignore polarization error as it is less than 0.001 radians (0.06%) [24, 25], which is negligible. We also ignore the Doppler effect [26, 27] as it can be compensated by precisely measuring the potential wavelength shift based on the trajectory and calibrating the photon wavelength accordingly [28]. Then, we combine all these phenomena to derive expressions for Fidelity and Efficiency of the protocol. Next, we compare these metrics against various important parameters, for which we use a realistic range of values. Lastly, we discuss our results and potential future directions towards the realization of uplink satellite networks. A detailed derivation of the Fidelity and Efficiency expressions can be found in the supplementary material, which, along with the Mathematica code containing the simulations, can be found here: [29].
@@ II The Protocol
Figure 1: (below) A schematic of the the proposed uplink setup. The setup consists of two ground stations sending one-half of their Bell pairs to a satellite via an uplink channel. The Bell measurement happens on the satellite, entangling the other two ground photons (represented by the dark blue curvy line). (above) The Bell-measurement apparatus inside the satellite. The apparatus consists of two optical modes, directed by a Polarizing Beam Splitter (PBS), followed by two 45 polarizers (one on each mode), and further followed by two PBSes, one on each mode, and foue bucket detectors, one on each output of the final PBSes.
@@ IX Results and discussion
Figure 2: Fidelity $F$ and success probability $\eta_{tot}$ vs time-gating window for various wavepacket widths $\sigma_{t}$ (in ns). The satellite altitude is taken to be 500km and the ground stations are separated by a distance of 1000km. All other parameters are chosen according to Table 1 of the Supplementary Material
@@
Results and Analysis. For our analysis, we consider the snapshot of the instant where the satellite is equidistant from the two ground stations. We are just considering the symmetry snapshot to simplify our calculations. In non-symmetric cases and cases that deviate from the symmetric snapshot, we assume that the satellite trajectory be pre-calculated and photon-release times and directions at the ground stations be appropriately adjusted to sync the photon arrival times at the satellite. This prevents any additional mode mismatch due to asymmetry. Therefore, all mode mismatch would be purely stemming from the limited precision in clock synchronization between the ground stations and the satellite.
To study how success probability (8) and fidelity (9) behave, we consider realistic parameter values chosen from various experiments and industry standards [see Table 1 and Section 9 of the supplementary material]. We organize our simulation around four main quantities of interest: temporal wavepacket width $\sigma_{t}$, time-gating window $t=t_{\mathrm{max}}-t_{\mathrm{min}}$, satellite altitude $h$ and distance $D_{G}$ between ground stations. The first two quantities have a large range of experimentally realizable values and affect the overall fidelity and $\eta_{tot}$ non-trivially, therefore their trends must be examined to establish a reasonable range of values for these parameters. Following this, we characterize the range of distances and satellite altitudes across which the protocol can produce feasible results.
@@
Best performance. Optimising the design of entanglement distribution in the framework of our protocol is a challenging task because the problem is a nonlinear optimization. Nevertheless, the most performant protocol we can reach is with a gating window of 40 ns, wavepacket width of 10ns, satellite altitude of 200km, and ground separation of 300km: this protocol distributes entangled state at a fidelity of 0.972 with an success probability of $1.5\times 10^{-4}$. Nonetheless, even with a satellite altitude of 500km and a ground separation as far as 1000km, we can obtain fidelity of 0.84 and a success probability of $2.404\times 10^{-6}$.